# On-Policy Policy Gradient Reinforcement Learning Without On-Policy Sampling

### v2

@@ 4 Correcting Sampling Error in Reinforcement Learning
$\nabla_{\bm{\theta}}J({\bm{\theta}})=\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\left[A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right].$ (4)
@@
The policy gradient is thus a linear combination of the gradient for each $({\bm{s}},{\bm{a}})$ pair $\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$ weighted by $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$. Let ${\mathcal{D}}$ be a dataset of trajectories. It is straightforward to show that the Monte Carlo estimate of the policy gradient can be written in a similar form as Equation 4 except with the true state-action visitation distribution replaced with the empirical visitation distribution, $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$ [Hanna et al., 2021]. Consequently, when $({\bm{s}},{\bm{a}})$ is over-sampled (i.e., $d_{\mathcal{D}}({\bm{s}},{\bm{a}})>d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$), the $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}|{\bm{s}})$ contributes more to the overall gradient than it should. Similarly, when $({\bm{s}},{\bm{a}})$ is under-sampled, $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}|{\bm{s}})$ contributes less than it should.
@@
We now provide a concrete example illustrating how small amounts of sampling error can cause the wrong actions to be reinforced, resulting in sub-optimal convergence. Suppose that in a particular state ${\bm{s}}_{0}$, an agent places equal probability on two actions ${\bm{a}}_{0}$ and ${\bm{a}}_{1}$ with advantages $A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}}_{0})=20$ and $A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}}_{1})=15$, respectively. Since $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{0}|{\bm{s}}_{0})=\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{1}|{\bm{s}}_{0})$ and $d_{{\pi_{\bm{\theta}}}}({\bm{s}}_{0},{\bm{a}}_{0})=d_{{\pi_{\bm{\theta}}}}({\bm{s}}_{0},{\bm{a}}_{1})$, the expected gradient will increase the probability of sampling the action with the larger advantage (${\bm{a}}_{0}$). With on-policy sampling, after 10 visits to ${\bm{s}}_{0}$, the agent will sample both actions 5 times in expectation. However, the agent may actually observe ${\bm{a}}_{0}$ 4 times and ${\bm{a}}_{1}$ 6 times. A Monte Carlo estimate of the policy gradient would then place $80.4\cdot A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}})=$ 8 weight on the gradient of ${\bm{a}}_{0}$ and $90.6\cdot A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}}_{1})=$ 9 weight on ${\bm{a}}_{1}$, thus decreasing the probability of sampling the optimal ${\bm{a}}_{0}$ action.
@@
Sampling error in on-policy sampling vanishes as the size of the batch of data used to estimate the gradient tends toward infinity. However, the preceding example suggests a simple strategy that would eliminate sampling error with finite data: have the agent adapt its probability on the next action it takes based on what actions it has already sampled. Continuing with our example, suppose the agent has visited ${\bm{s}}_{0}$ 9 times and sampled ${\bm{a}}_{0}$ 4 times and ${\bm{a}}_{1}$ 5 times. With on-policy sampling, the agent may observe ${\bm{a}}_{1}$ again upon the next visit to ${\bm{s}}_{0}$. Alternatively, the agent could sample its next action from a distribution that puts probability 1 on ${\bm{a}}_{0}$ and consequently produce an aggregate batch of data that contains both actions in their expected frequency. While this adaptive method is an off-policy sampling method, it produces data that exactly matches the on-policy distribution and will thus produce a more accurate gradient.
@@
This example suggests that we can heuristically reduce sampling error by taking the most under-sampled action at a given state. Under a strong assumption that the MDP had a DAG structure, Zhong et al. [2022] proved that this heuristic results in the empirical distribution of states and actions in a fixed-horizon MDP converging to $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$ and moreover converging at a faster rate than on-policy sampling. We remove the limiting DAG assumption with the following result:
@@
Assume that data is collected with an adaptive behavior policy that always takes the most under-sampled action in each state, $s$, with respect to policy $\pi$, i.e, $a\leftarrow\arg\max_{a^{\prime}}(\pi(a^{\prime}|s)-{\pi_{\mathcal{D}}}(a^{\prime}|s))$, where ${\pi_{\mathcal{D}}}$ is the empirical policy after $m$ state-action pairs have been collected. Assume that ${\mathcal{S}}$ and ${\mathcal{A}}$ are finite and that the Markov chain induced by $\pi$ is irreducible. Then we have that the empirical state visitation distribution, $d_{m}$, converges to the state distribution of $\pi$, $d_{\pi}$, with probability $1$:
@@
Proof.
See Appendix A. ∎
While adaptively sampling the most under-sampled action can reduce sampling error, this heuristic is difficult to implement in practice. In tasks with continuous states and actions, the $\operatorname*{arg\,max}$ often has no closed-form solution, and the empirical policy can be expensive to compute at every timestep. Building upon the concepts discussed in this section, the following section presents a scalable adaptive sampling algorithm that reduces sampling error in on-policy policy gradient learning.
@@ 5 Proximal Robust On-Policy Sampling for Policy Gradient Algorithms
Our goal is to develop an adaptive, off-policy sampling algorithm that reduces sampling error in on-policy data collection for on-policy policy gradient algorithms. We outline a general framework for on-policy learning with an adaptive behavior policy in Algorithm 1. In this framework, the behavior policy $\pi_{\bm{\phi}}$ and target policy $\pi_{\bm{\theta}}$ are initially the same. The behavior policy collects a batch of $m$ transitions, adds the batch to a data buffer ${\mathcal{D}}$, and then updates its weights such that the next batch it collects reduces sampling error in ${\mathcal{D}}$ with respect to the target policy $\pi_{\bm{\theta}}$ (Lines 7-10). Every $n$ steps (with $n>m$), the agent updates its target policy with data from ${\mathcal{D}}$ (Line  11). We refer to $m$ and $n$ as the behavior batch size and the target batch size, respectively.
@@
1:  Inputs: Target batch size $n$, behavior batch size $m$, buffer size $b$.
@@
5:  Initialize empty buffer ${\mathcal{D}}$ with capacity $bn$.
@@
A subtle implication of adaptive sampling is that it can correct sampling error in any empirical data distribution – even one generated by a different policy. Rather than discarding off-policy data from old policies – as is commonly done in on-policy learning – we let the data buffer hold up to $b$ target batches ($bn$ transitions) and call $b$ the buffer size. If $b>1$, then ${\mathcal{D}}$ will contain historic off-policy data used in previous target policy updates. Regardless of how $b$ is set, the role of the behavior policy is to continually adjust action probabilities for new samples so that the aggregate data distribution of ${\mathcal{D}}$ matches the expected on-policy distribution of the current target policy (Line 10). Implementing Line 10 is the core challenge we address in the remainder of this section.
To ensure that the empirical distribution of ${\mathcal{D}}$ matches the expected on-policy distribution, updates to $\pi_{\bm{\phi}}$ should attempt to increase the probability of actions which are currently under-sampled with respect to $\pi_{\bm{\theta}}$. Zhong et al. [2022] recently developed a method called Robust On-policy Sampling (ros) for making such updates. In particular, the gradient $\nabla_{\bm{\phi}}{\mathcal{L}}:=-\nabla_{\bm{\phi}}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ when evaluated at ${\bm{\phi}}={\bm{\theta}}$ provides a direction to change ${\bm{\phi}}$ such that under-sampled actions have their probabilities increased. Thus a single step of gradient ascent will increase the probability of under-sampled actions. [^1] To add further intuition for this update, note that it is the opposite of a gradient ascent step on the log likelihood of ${\mathcal{D}}$. When starting at ${\bm{\theta}}$, gradient ascent on the data log likelihood will increase the probability of actions that are over-sampled relative to $\pi_{\bm{\theta}}$. Hence, the ros update changes ${\bm{\phi}}$ in the opposite direction. In theory and in simple rl policy evaluation tasks, this update was shown to improve the rate at which the empirical data distribution converges to the on-policy distribution – even when the empirical data distribution contains off-policy data. Unfortunately, there are two main challenges that render ros unsuitable for Line 10 in Algorithm 1.
@@
1:  Inputs: Target policy parameters ${\bm{\theta}}$, buffer ${\mathcal{D}}$, target kl $\delta$, clipping coefficient $epsilon_{\textsc{props}}{}{}$, regularizer coefficient $\lambda$, n_epoch, n_minibatch.
@@
7:        Compute the loss (Eq. 6)
${\mathcal{L}}\leftarrow\frac{1}{|{\mathcal{D}}_{j}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}_{j}}{\mathcal{L}}_{\textsc{props}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\textsc{props}}{}{},\lambda)$
8:        Update ${\bm{\phi}}$ with a step of gradient ascent on ${\mathcal{L}}$
9:        if $D_{\textsc{kl}{}}(\pi_{\bm{\theta}}||\pi_{\bm{\phi}})>\delta_{\textsc{props}}{}$ then
@@
Challenge 1: Destructively large policy updates. Since the buffer ${\mathcal{D}}$ may contain data collected from older target policies, some samples in ${\mathcal{D}}$ may be very off-policy with respect to the current target policy such that $\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ is large and negative. Since $\nabla_{\bm{\phi}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ increases in magnitude as $\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ tends towards zero, ros incentivizes the agent to continually decrease the probability of these actions despite being extremely unlikely under the current target policy. Thus, off-policy samples can produce destructively large policy updates.
[…]

### v3

@@ 4 Correcting Sampling Error in Reinforcement Learning
$\nabla_{\bm{\theta}}J({\bm{\theta}})=\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\left[A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right].$ (4)
@@
The policy gradient is thus a linear combination of the gradient for each $({\bm{s}},{\bm{a}})$ pair $\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$ weighted by $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$. Let $\mathcal{D}$} be a dataset of trajectories. It is straightforward to show that the Monte Carlo estimate of the policy gradient can be written in a form similar to Eq. 4 except with the true state-action visitation distribution replaced with the empirical visitation distribution $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$, denoting the fraction of times $({\bm{s}},{\bm{a}})$ appears in $\mathcal{D}$} (Hanna et al., 2021):
$\nabla_{\bm{\theta}}\widehat{J}({\bm{\theta}})=\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}d_{\mathcal{D}}({\bm{s}},{\bm{a}})\left[A^{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right].$ (5)
By comparing Eq. 4 and Eq. 5, we can understand how sampling error affects the Monte Carlo policy gradient estimate. When $({\bm{s}},{\bm{a}})$ is over-sampled (i.e., $d_{\mathcal{D}}({\bm{s}},{\bm{a}})>d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$), then $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}|{\bm{s}})$ contributes more to the overall gradient than it should. Similarly, when $({\bm{s}},{\bm{a}})$ is under-sampled, $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}|{\bm{s}})$ contributes less than it should. Below, we provide a concrete example illustrating how small amounts of sampling error can cause the wrong actions to be reinforced.
@@
Example: Sampling error can cause incorrect policy updates. Let $\pi_{\bm{\theta}}$ be a policy for an MDP with two discrete actions ${\bm{a}}_{0}$ and ${\bm{a}}_{1}$, and suppose that in a particular state ${\bm{s}}_{0}$, the advantage of each action w.r.t. $\pi_{\bm{\theta}}$} is $A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}}_{0})=20$ and $A^{\pi_{\bm{\theta}}}({\bm{s}}_{0},{\bm{a}}_{1})=15$. For simplicity, suppose the policy has a direct parameterization $\pi_{\bm{\theta}}({\bm{a}}_{0}|{\bm{s}})={\bm{\theta}}_{{\bm{s}}}$, $\pi_{\bm{\theta}}({\bm{a}}_{1}|{\bm{s}})=1-{\bm{\theta}}_{{\bm{s}}}$ and places equal probability on both actions in ${\bm{s}}_{0}$ (${\bm{\theta}}_{{\bm{s}}_{0}}=0.5)$. Then, we have $\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{0}|{\bm{s}}_{0})=-\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{1}|{\bm{s}}_{0})$ and $d_{{\pi_{\bm{\theta}}}}({\bm{s}}_{0},{\bm{a}}_{0})=d_{{\pi_{\bm{\theta}}}}({\bm{s}}_{0},{\bm{a}}_{1})$ so that the expected gradient increases the probability of sampling ${\bm{a}}_{0}$, the optimal action. With on-policy sampling, after 10 visits to ${\bm{s}}_{0}$, the agent will sample both actions 5 times in expectation. However, if the agent actually observes ${\bm{a}}_{0}$ 4 times and ${\bm{a}}_{1}$ 6 times, a Monte Carlo estimate of the policy gradient then yields $\frac{4}{10}\cdot 20\cdot\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{0}|{\bm{s}}_{0})+\frac{6}{10}\cdot 15\cdot\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{1}|{\bm{s}}_{0})=-\nabla_{\bm{\theta}}\log{\pi_{\bm{\theta}}}({\bm{a}}_{0}|{\bm{s}}_{0})$ which decreases the probability of sampling the optimal ${\bm{a}}_{0}$ action.
@@
Sampling error in on-policy sampling vanishes as the size of the batch of data used to estimate the gradient tends toward infinity. However, the preceding example suggests a simple strategy to eliminate sampling error with finite data: have the agent adapt its probability on the next action it takes based on the actions it has already sampled. Continuing with our example, suppose the agent has visited ${\bm{s}}_{0}$ 9 times and sampled ${\bm{a}}_{0}$ 4 times and ${\bm{a}}_{1}$ 5 times. With on-policy sampling, the agent may observe ${\bm{a}}_{1}$ again upon the next visit to ${\bm{s}}_{0}$. Alternatively, the agent could sample its next action from a distribution that puts probability 1 on ${\bm{a}}_{0}$ and consequently produce an aggregate batch of data that contains both actions in their expected frequency. While this adaptive method uses off-policy sampling, it produces data that exactly matches the on-policy distribution and thus produces a more accurate gradient.
@@
This example suggests that we can heuristically reduce sampling error by selecting the most under-sampled action at a given state. Under a strong assumption that the MDP has a directed acyclic graph (DAG) structure, Zhong et al. (2022) proved that in a fixed-horizon MDP, this heuristic produces an empirical state-action distribution that converges to $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$ and moreover converges at a faster rate than on-policy sampling. We remove this limiting DAG assumption with the following result:
@@
Assume that data is collected with an adaptive behavior policy that always takes the most under-sampled action in each state $s$ w.r.t. $\pi$, i.e., $a\leftarrow\arg\max_{a^{\prime}}(\pi(a^{\prime}|s)-{\pi_{\mathcal{D}}}(a^{\prime}|s))$, where $\pi_{\mathcal{D}}$} is the empirical policy after $m$ state-action pairs have been collected. Assume that $\mathcal{S}$} and $\mathcal{A}$} are finite and that the Markov chain induced by $\pi$ is irreducible. Then we have that the empirical state visitation distribution, $d_{m}$, converges to the state distribution of $\pi$, $d_{\pi}$, with probability $1$:
@@
We prove Proposition 1 in Appendix A. While adaptively sampling the most under-sampled action can reduce sampling error, this heuristic is difficult to implement in practice; in tasks with continuous states and actions, the $\operatorname*{arg\,max}$ in Proposition 1 often has no closed-form solution, and the empirical policy $\pi_{\mathcal{D}}$ can be expensive to compute at every timestep. Building upon the concepts discussed in this section, the following section presents a scalable adaptive sampling algorithm that reduces sampling error in on-policy policy gradient learning.
@@ 5 Proximal Robust On-Policy Sampling for Policy Gradient Algorithms
1:Inputs: Target batch size $n$, behavior update frequency $m$, buffer size $b$.
@@
5: Initialize empty buffer $\mathcal{D}$} with capacity bnbn.
@@
Our goal is to develop an adaptive, off-policy sampling algorithm that reduces sampling error in on-policy data collection for on-policy policy gradient algorithms. We outline a general framework for on-policy learning with an adaptive behavior policy in Algorithm 1. The behavior policy collects a batch of $m$ transitions, adds the batch to a data buffer $\mathcal{D}$}, and then updates its weights such that the next batch it collects reduces sampling error in $\mathcal{D}$} with respect to the target policy $\pi_{\bm{\theta}}$ (Lines 7-10). Every $n$ steps (with $n>m$), the agent updates its target policy with data from $\mathcal{D}$} (Line  11). We refer to $m$ and $n$ as the behavior update frequency and the target batch size, respectively.
@@
The behavior policy must continually adjust action probabilities for new samples so that the aggregate data distribution of $\mathcal{D}$} matches the expected on-policy distribution of the current target policy (Line 10). A subtle implication of this adaptive sampling is that it can correct sampling error in any empirical data distribution—even one generated by a different policy. Thus, rather than discarding off-policy data from old policies as is commonly done in on-policy learning, we let the data buffer hold up to $b$ target batches ($bn$ transitions) and call $b$ the buffer size. If $b>1$, then $\mathcal{D}$} will contain historic off-policy data used in previous target policy updates. [^3] The advantage estimates in historic data correspond to historic policies and are thus biased advantage estimates of the current policy. In Section 7, we provide additional detail on how to mitigate this bias with generalized advantage estimation (GAE). Implementing Line 10 is the core challenge we address in the remainder of this section.
5.1 Robust On-Policy Sampling
To ensure that the empirical distribution of $\mathcal{D}$} matches the expected on-policy distribution, updates to $\pi_{\bm{\phi}}$ should attempt to increase the probability of actions which are currently under-sampled with respect to $\pi_{\bm{\theta}}$. Zhong et al. (2022) recently developed a simple method called Robust On-policy Sampling (ROS) for making such updates, which PROPS builds upon. In this section, we briefly review ROS and discuss its limitations.
Let $\pi_{\mathcal{D}}({\bm{a}}|{\bm{s}})$ denote the empirical policy, the fraction of times action $\bm{a}$} is taken at state $\bm{s}$} in $\mathcal{D}$}. If $\pi_{\mathcal{D}}({\bm{a}}|{\bm{s}})<\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$, then $\bm{a}$} appears fewer times at $\bm{s}$} than it would in expectation under $\pi_{\bm{\theta}}$ and is therefore under-sampled. Thus, $\pi_{\mathcal{D}}$ provides a way to identify actions whose probabilities should be increased to reduce sampling error. In the tabular setting, $\pi_{\mathcal{D}}$ can be computed exactly as the maximizer of the log-likelihood,
$\mathcal{L}({\bm{\theta}}^{\prime})=\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\log\pi_{{\bm{\theta}}^{\prime}}({\bm{a}}\mid{\bm{s}}),$ (6)
i.e., $\pi_{\mathcal{D}}=\arg\max_{\bm{\phi}}\mathcal{L}({\bm{\theta}}^{\prime})$. However, in tasks with high-dimensional continuous state and action spaces, computing $\pi_{\mathcal{D}}$ explicitly is generally intractable. The key insight of ROS is that we do not need to compute $\pi_{\mathcal{D}}$ explicitly; we only need a direction that tells us how to adjust action probabilities based on how the data in $\mathcal{D}$} differs from $\pi_{\theta}$. The gradient of the log-likelihood at ${\bm{\phi}}={\bm{\theta}}$, $\nabla_{\bm{\phi}}\mathcal{L}_{(}{\bm{\phi}})\big|_{{\bm{\phi}}={\bm{\theta}}}$, points in a direction that moves $\pi_{\bm{\phi}}$ toward $\pi_{\mathcal{D}}$, so a step in this direction decreases the probabilities of actions that are under-sampled in $\mathcal{D}$} w.r.t. $\pi_{\bm{\theta}}$ and increases those of over-sampled actions. Therefore, taking a step in the opposite direction increases the probabilities of under-sampled actions and decreases those of over-sampled actions, providing a way to reduce sampling error without ever computing $\pi_{\mathcal{D}}$. Thus, ROS performs a single step of descent on the negative log-likelihood
$\nabla_{\bm{\phi}}\mathcal{L}({\bm{\phi}})\big|_{{\bm{\phi}}={\bm{\theta}}}=\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}-\nabla_{\bm{\phi}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})|_{{\bm{\phi}}={\bm{\theta}}},$ (7)
at each timestep to increase the probability of under-sampled actions. In theory and in low-dimensional RL policy evaluation tasks such as CartPole-v1 (Brockman et al., 2016), this update was shown to improve the rate at which the empirical data distribution converges to the on-policy distribution—even when the empirical data distribution contains off-policy data. Unfortunately, two main challenges render ROS unsuitable for Line 10 in Algorithm 1.
@@ 5.1 Robust On-Policy Sampling
1:Inputs: Target policy parameters $\bm{\theta}$}, buffer $\mathcal{D}$}, target KL $\delta_{\text{PROPS}}$, clipping coefficient $\epsilon_{\text{PROPS}}{}$, regularizer coefficient $\lambda$, n_epoch, n_minibatch.
@@
7:   Update $\bm{\phi}$} with a step of gradient ascent on loss
$\frac{1}{|{\mathcal{D}}_{j}|}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}_{j}}{\mathcal{L}}_{\text{PROPS}}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\text{PROPS}},\lambda)$
8:   if $D_{\text{KL}}(\pi_{\bm{\theta}}||\pi_{\bm{\phi}})>\delta_{\text{PROPS}}$ then
9:    return $\bm{\phi}$}
@@
Challenge 1: Destructively large policy updates. Since $\mathcal{D}$} may contain data collected by old target policies, some samples in $\mathcal{D}$} may be very off-policy w.r.t. the current target policy such that $\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ is large and negative. Since $\nabla_{\bm{\phi}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ increases in magnitude as $\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$ approaches zero, those off-policy samples can produce destructively large updates.
Challenge 2: Improper handling of continuous actions. In continuous-action tasks, ROS may produce behavior policies that increase sampling error. Continuous policies $\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})$ are typically parameterized as Gaussians ${\mathcal{N}}({\bm{\mu}}({\bm{s}}),\Sigma({\bm{s}}))$ with mean ${\bm{\mu}}({\bm{s}})$ and diagonal covariance matrix $\Sigma({\bm{s}})$. Since actions in the tail of the Gaussian far from the mean will generally be under-sampled, the ROS update will continually push the components of ${\bm{\mu}}({\bm{s}})$ towards $\pm\infty$ and the diagonal components of $\Sigma({\bm{s}})$ towards 0 to increase the probability of sampling these actions. The result is a degenerate behavior policy that is so far from the target policy that sampling from it increases sampling error. We illustrate this scenario with 1-dimensional continuous actions in Fig. 7 of Appendix C. In the next section, we discuss how to address both of these challenges.
5.2 Proximal Robust On-Policy Sampling